OpenStudy (anonymous):

If X is continuous uniform on the interval (0, 10), find P (X+ (10/X) > 7). @Statistics

OpenStudy (anonymous):

normally you cannot multiply both sides of an inequality by x, but you can in this case because you know that x is positive.

OpenStudy (anonymous):

so you get \[x+\frac{10}{x}>7\] \[x^2+10>7x\] \[x^2-7x+10>\] \[(x-2)(x-5)>0\] and so \[x<2,x>5\]

OpenStudy (anonymous):

therefore the lengths are from 0 to 2 is 2 and from 5 to 10 is 5 for a total length of 7 divide by the length of the entire interlval to get \[P(X+\frac{10}{X}>7)=\frac{7}{10}=.7\]

OpenStudy (anonymous):

thanks man! can youi help me on my other question, please?